OpenStudy (anonymous):

Let f(x) = -2x - 7 and g(x) = -4x + 6. find g of f (-5)

OpenStudy (anonymous):

You can do this one of two methods: 1. Find f(-5) then find g of that number 2. Find an equation for g(f(x)) and put -5 in 1. f(-5)=-2(-5)-7=3 g(3)=-4(3)+6=-6 2. g(f(x))=-4(-2x-7)+6=8x+34 g(f(-5))=8(-5)+34=-6 Both result in the same answer, take your pick :)

OpenStudy (anonymous):

still confused

OpenStudy (anonymous):

Do you understand how a function works? You put one number in and get a new number out. g(f(x)) means you put x through f to get a new number out, then you put that new number through g.

OpenStudy (anonymous):

i got it

OpenStudy (anonymous):

if f(x)=3x+2, what is the equation for f–1(x)?